OpenStudy (surana):

Solve. (x-3)^2 = 25.

OpenStudy (surana):

The ^2 represents exponent 2.

OpenStudy (surana):

@millsemily Just wanted to see if you could help.

OpenStudy (welshfella):

you need to take the square root of both sides

OpenStudy (welshfella):

the expression (x -3)^2 is a perfect square so what is its square root?

OpenStudy (welshfella):

^2 means 'squared'

OpenStudy (michele_laino):

hint: \[{\left( {x - 3} \right)^2} - 25 = {\left( {x - 3} \right)^2} - {5^2} = \left( {x - 3 - 5} \right)\left( {x - 3 + 5} \right) = \left( {x - 8} \right)\left( {x + 2} \right)\]

OpenStudy (anonymous):

what number multiplied by itself =25 ? this is what you think when asked for the square root of 25 so you know it is 5*5 so that's the right side solved=5 the left is what bracket multiplied by itself =(x-3)^2 it is (x-3)(x-3) so square root is (x-3)

OpenStudy (welshfella):

Hint: there are 2 square roots of 25

OpenStudy (surana):

So the calculations made mean the answer is X = +/- 3? Because that's what I came up with if I follow the calculations here.

OpenStudy (surana):

I think it's 8 and -2, given what I got when I rewrote it.

OpenStudy (surana):

Thanks everyone.